# Commutator calculus and symbolic differentiation of matrix functions

### v2

@@ 3.2. Further identities for the derivative of the matrix power and logarithm
With the help of the calculus rules in Theorem 2.2, we can now apply (3.57) and (3.58) in (3.51) and (3.52), leading to (3.53) and (3.54), respectively. This also readily proves the reverse implications in (3.55) and (3.56). Conversely, assume that $\mu(\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}})\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}}^{2}\mathbf{X}={\boldsymbol{0}}$. By the positivity of $\mu$ and Theorem 2.2 (iv), this is equivalent to $\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}}^{2}\mathbf{X}={\boldsymbol{0}}$. However, due to Theorem 2.2 (vi), this implies that $\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}}\mathbf{X}={\boldsymbol{0}}$ since
@@ 3.3. Possible extension to non-symmetric matrices
For the definition of a logarithm of a general matrix, we can refer to [36, Th. 1]. Then, there appear to be numerous approaches towards defining $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$ for $\mathbf{G}\in\mathbb{C}^{d\times d}$. First, of course, one can apply the general theory of holomorphic calculus to the operator $\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$ (and similar). However, this requires that the applied function is holomorphic in an open subset of $\mathbb{C}$ that contains the spectrum of the operator, even in the case where the spectrum is actually real. The second possibility is to replace the Schur decomposition of $\mathbf{G}$ by the singular value decomposition. This obviously places the least requirements on the applied functions; on the other hand, one has to consider an appropriate notion of a generalized matrix function and its derivative, leading to more involved formulae than those in (2.2) and (3.3), see [15] and especially [27] for details. Another approach is to use the spectral measure, that also offers a way to extend the theory to general self-adjoint operators. For this, see the original Russian literature, cf. [9], [11], [12] and references therein. Further alternative approaches to defining and evaluating matrix functions and their derivatives (e.g. using polynomial interpolation), can be found in the works by Higham & Al-Mohy (cf. [1], [17]). As we are here primarily concerned with applications of the commutator calculus to symmetric matrices, we avoid any of the above technicalities.
@@ 4.1. Monotonicity of matrix functions
For $f\in\mathcal{C}^{1}(I;\mathbb{R})$, we can apply the mean value theorem to the function $s\mapsto f(\mathbf{H}+s(\mathbf{G}-\mathbf{H}))\cdot(\mathbf{G}-\mathbf{H})$ and obtain this way $\mathbf{G}_{\xi}:=\mathbf{H}+\xi(\mathbf{G}-\mathbf{H})$, $\xi\in(0,1)$, such that
@@
$\big{(}f(\mathbf{G})-f(\mathbf{H})\big{)}\cdot(\mathbf{G}-\mathbf{H})=\Bigg{|}\sqrt{\frac{f(\operatorname{\mathsf{L}}_{\mathbf{G}_{\xi}})-f(\operatorname{\mathsf{R}}_{\mathbf{G}_{\xi}})}{\operatorname{\mathsf{L}}_{\mathbf{G}_{\xi}}-\operatorname{\mathsf{R}}_{\mathbf{G}_{\xi}}}}(\mathbf{G}-\mathbf{H})\Bigg{|}^{2},$
@@
This agrees with the results of [20, Prop. 4.8] obtained by different methods. It is also quite obvious here that the assumption of continuous differentiability of $f$ could be relaxed if needed.
@@ 4.4. Logarithmic reformulation of the classical viscoelastic fluid models
${\boldsymbol{\Omega}}^{\log}:=\mathbf{W}-\sigma(2\operatorname{\mathsf{a\!d}}_{\mathbf{H}})\mathbf{D}$ (4.13)
@@
$\sigma(x):=\Big{\{}\begin{matrix}\coth x-\frac{1}{x}&x\neq 0,\\ 0&x=0.\end{matrix}$ (4.14)
Note that the function $\sigma$ is continuous and odd, hence the tensor ${\boldsymbol{\Omega}}^{\log}$ is well defined by (4.13) (appealing to Definition 2.1) and antisymmetric due to Theorem 2.2 (v).
@@
$\boldsymbol{\Omega}=\mathbf{W}-f(\operatorname{\mathsf{a\!d}}_{\mathbf{H}})\mathbf{D},\qquad f\text{ continuous and odd.}$. (4.15)
@@ 4.5. Connecting the logarithmic and upper-convected rates
Analogous relations to (4.22) and (4.23) can, of course, be deduced for other types of objective derivative used in various physics applications; see [26, (1.13), (1.16)] for further examples. We also refer to [3], [13], [26], [32] and references therein for the state-of-the-art developments and applications of objective rates and logarithmic strain in continuum mechanics.
@@ 4.6. Comparing dissipation potentials
When equation (4.18) is supplied with the diffusion term $-\Delta\mathbf{B}$, it corresponds to the additional term in the dissipation of the form
@@
indicating that the simplest diffusion term in the formulation using $\mathbf{B}$ has generally a stronger dissipative effect than its counterpart in the logarithmic formulation using $\mathbf{H}$.
@@ 5. Proofs
We recall that $[m_{ij}]$ denotes a $d\times d$ matrix with elements $m_{ij}$, $i,j=1,\ldots,d$.
@@ 5.1. Functions applied to the commutator
the linear problem (2.8) admits a unique solution if all the eigenvalues of $\mathbf{G}$ are distinct. In case of eigenvalue multiplicities, duplicate entries can be eliminated in $\mathbf{V}_{\mathbf{G}}$ and $[f(g_{i},g_{j})]$, leading to an underdetermined problem for $\mathbf{J}_{\mathbf{G}}^{f}$ that can be solved using generalized matrix inverse methods.
@@
$(\mathbf{G}^{m}\mathbf{X}\mathbf{G}^{n})_{ij}=g_{i}^{m}g_{j}^{n}\mathbf{X}_{ij}.$
@@
$(\mathbf{G}\mathbf{X}-\mathbf{X}\mathbf{G})_{ij}=(g_{i}-g_{j})\mathbf{X}_{ij}=(g_{1}-g_{2})\left[\begin{matrix}0&1\\ -1&0\end{matrix}\right]_{ij}\mathbf{X}_{ij}.$ (5.5)
Using the last two relations, we find
@@
for any $k\in\{1,2,3\}$. Now we are ready to evaluate the terms on the right-hand side of (2.13). For the part depending on $f_{\rm odd}$, we have
@@
$\qquad=\left[\begin{matrix}0&f_{\rm even}(\frac{g_{1}-g_{2}}{2})&f_{\rm even}(\frac{g_{3}-g_{1}}{2})\\ f_{\rm even}(\frac{g_{1}-g_{2}}{2})&0&f_{\rm even}(\frac{g_{2}-g_{3}}{2})\\ f_{\rm even}(\frac{g_{3}-g_{1}}{2})&f_{\rm even}(\frac{g_{2}-g_{3}}{2})&0\end{matrix}\right]_{ij}\mathbf{X}_{ij}.$
Hence, by summing the last two identities and adding also the part $f_{0}\mathbf{X}$, we arrive at (2.13).
Regarding the eigenvalue multiplicities, the case of a single distinct eigenvalue always trivially leads to (2.11). Next, suppose that $d=3$ and $g_{1}\neq g_{2}=g_{3}$. We proceed just as in (5.4) and (5.5) with the difference that $\mathbf{G}$ and $\mathbf{X}$ are now of the size $\times 3$ 3. We also note that for the additional elements, we have
@@
which again leads to the required result. The other two cases $g_{2}\neq g_{3}=g_{1}$ and $g_{3}\neq g_{1}=g_{2}$ are analogous. ∎
@@ 5.2. Symbolic differentiation of matrix functions
[…]
@@
[…]
@@
(B) Using the well-known identity for the beta function ([30, Sect. 6.2])
[…]
@@
[…]
@@
[…]
@@
$\frac{d\log\mathbf{A}}{d\mathbf{A}}=\frac{\log\operatorname{\mathsf{L}}_{\mathbf{A}}-\log\operatorname{\mathsf{R}}_{\mathbf{A}}}{\operatorname{\mathsf{L}}_{\mathbf{A}}-\operatorname{\mathsf{R}}_{\mathbf{A}}}=\frac{\operatorname{\mathsf{L}}_{\log\mathbf{A}}-\operatorname{\mathsf{R}}_{\log\mathbf{A}}}{e^{\operatorname{\mathsf{L}}_{\log\mathbf{A}}}-e^{\operatorname{\mathsf{R}}_{\log\mathbf{A}}}}$ $=e^{-\operatorname{\mathsf{R}}_{\log\mathbf{A}}}\frac{\operatorname{\mathsf{L}}_{\log\mathbf{A}}-\operatorname{\mathsf{R}}_{\log\mathbf{A}}}{e^{\operatorname{\mathsf{L}}_{\log\mathbf{A}}-\operatorname{\mathsf{R}}_{\log\mathbf{A}}}-1}=\operatorname{\mathsf{R}}_{\mathbf{A}^{-1}}(1/\eta)(2\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}}),$
@@
hence, by integration of this identity over (0,1)(0,1)( 0 , 1 ), we get
@@
$\frac{\sinh(rx)}{\sinh x\cosh((r-1)x)}$ $=\frac{\sinh x\cosh((r-1)x)+\cosh x\sinh((r-1)x)}{\sinh x\cosh((r-1)x)}=1+\frac{\tanh((r-1)x)}{\tanh{x}},$

### v3

@@ 3.2. Further identities for the derivative of the matrix power and logarithm
With the help of the calculus rules in Theorem 2.2, we can now apply (3.57) and (3.58) in (3.51) and (3.52), leading to (3.53) and (3.54), respectively. This also readily proves the reverse implications in (3.55) and (3.56). Conversely, assume that $\mu(\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}})\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}}^{2}\mathbf{X}={\boldsymbol{0}}$. By the positivity of $\mu$ and Theorem 2.2 (iv), this is equivalent to $\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}}^{2}\mathbf{X}={\boldsymbol{0}}$. However, due to the symmetry of $\mathbf{A}$ and using Theorem 2.2 (vi), this actually implies that $\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}}\mathbf{X}={\boldsymbol{0}}$ since
@@ 3.3. Possible extension to non-symmetric matrices
For the definition of a logarithm of a general matrix, we can refer to [37, Th. 1]. Then, there appear to be numerous approaches towards defining $f(\operatorname{\mathsf{a\!d}}_{\mathbf{G}})$ for $\mathbf{G}\in\mathbb{C}^{d\times d}$. First, of course, one can apply the general theory of holomorphic calculus to the operator $\operatorname{\mathsf{a\!d}}_{\mathbf{G}}$ (and similar). However, this requires that the applied function is holomorphic in an open subset of $\mathbb{C}$ that contains the spectrum of the operator, even in the case where the spectrum is actually real. The second possibility is to replace the Schur decomposition of $\mathbf{G}$ by the singular value decomposition. This obviously places the least requirements on the applied functions; on the other hand, one has to consider an appropriate notion of a generalized matrix function and its derivative, leading to more involved formulae than those in (2.2) and (3.3), see [15] and especially [28] for details. Another approach is to use the spectral measure, that also offers a way to extend the theory to general self-adjoint operators. For this, see the original Russian literature, cf. [9], [11], [12] and references therein. Further alternative approaches to defining and evaluating matrix functions and their derivatives (e.g. using polynomial interpolation), can be found in the works by Higham & Al-Mohy (cf. [1], [17]). As we are here primarily concerned with applications of the commutator calculus to symmetric matrices, we avoid any of the above technicalities.
@@ 4.1. Monotonicity of matrix functions
Using polynomial interpolation methods, it can be shown that if $f\in\mathcal{C}^{1}(I;\mathbb{R})$, then the function
$g(s):=f(\mathbf{H}+s(\mathbf{G}-\mathbf{H}))\cdot(\mathbf{G}-\mathbf{H})$
inherits the property $g\in\mathcal{C}^{1}(I;\mathbb{R})$, see [20, Th. 6.6.14]. Hence, one can apply the mean value theorem to $g$ and obtain this way $\mathbf{G}_{\xi}:=\mathbf{H}+\xi(\mathbf{G}-\mathbf{H})$, $\xi\in(0,1)$, such that
@@
$\big(f(\mathbf{G})-f(\mathbf{H})\big)\cdot(\mathbf{G}-\mathbf{H})=\Bigg|\sqrt{\frac{f(\operatorname{\mathsf{L}}_{\mathbf{G}_{\xi}})-f(\operatorname{\mathsf{R}}_{\mathbf{G}_{\xi}})}{\operatorname{\mathsf{L}}_{\mathbf{G}_{\xi}}-\operatorname{\mathsf{R}}_{\mathbf{G}_{\xi}}}}(\mathbf{G}-\mathbf{H})\Bigg|^{2},$ (4.1)
@@
This agrees with the results of [21, Prop. 4.8] obtained by different methods.
@@ 4.4. Logarithmic reformulation of the classical viscoelastic fluid models
${\boldsymbol{\Omega}}^{\log}:=\mathbf{W}-\mathcal{L}(2\operatorname{\mathsf{a\!d}}_{\mathbf{H}})\mathbf{D}$ (4.14)
@@
$\mathcal{L}(x):=\Big\{\begin{matrix}\coth x-\frac{1}{x}&x\neq 0,\\ 0&x=0\end{matrix}$ (4.15)
is the Langevin function. Note that $\mathcal{L}$ is continuous and odd, hence the tensor ${\boldsymbol{\Omega}}^{\log}$ is well defined by (4.14) (appealing to Definition 2.1) and antisymmetric due to Theorem 2.2 (v).
@@
$\boldsymbol{\Omega}=\mathbf{W}-f(\operatorname{\mathsf{a\!d}}_{\mathbf{H}})\mathbf{D},\qquad f\text{ continuous and odd.}$ (4.16)
@@ 4.5. Connecting the logarithmic and upper-convected rates
Analogous relations to (4.23) and (4.24) can, of course, be deduced for other types of objective derivative used in various physics applications; see [27, (1.13), (1.16)] for further examples. We also refer to [3], [13], [27], [33] and references therein for the state-of-the-art developments and applications of objective rates and logarithmic strain in continuum mechanics.
@@ 4.6. Comparing dissipation potentials
When equation (4.19) is supplied with the diffusion term $\Delta\mathbf{B}$, it corresponds to the additional term in the dissipation of the form
@@
indicating that the simplest diffusion term in the formulation with $\mathbf{B}$ has generally a stronger dissipative effect than its counterpart in the logarithmic formulation using $\mathbf{H}$.
@@ 5.1. Functions applied to the commutator
We remind the reader that by $m_{ij}]$, we denote a $d\times d$ matrix with elements $m_{ij}$, $i,j=1,\ldots,d$.
@@
the linear problem (2.8) admits a unique solution if all the eigenvalues of $\mathbf{G}$ are distinct. In case of eigenvalue multiplicities, duplicate entries can be eliminated in $\mathbf{V}_{\mathbf{G}}$ and $f(g_{i},g_{j})]$, leading to an underdetermined problem for $\mathbf{J}_{\mathbf{G}}^{f}$ that can be solved using generalized matrix inverse methods.
@@
$(\mathbf{G}^{m}\mathbf{X}\mathbf{G}^{n})_{ij}=g_{i}^{m}g_{j}^{n}\mathbf{X}_{ij},\quad i,j=1,\ldots,d.$
@@
$(\mathbf{G}\mathbf{X}-\mathbf{X}\mathbf{G})_{ij}=(g_{i}-g_{j})\mathbf{X}_{ij}=(g_{1}-g_{2})\left[\begin{matrix}0&1\\ -1&0\end{matrix}\right]_{ij}\mathbf{X}_{ij}$ (5.5)
for each $i,j\in\{1,2\}$. Using the last two relations, we find
@@
for any $i,j,k\in\{1,2,3\}$. Now we are ready to evaluate the terms on the right-hand side of (2.13). For the part depending on $f_{\rm odd}$, we have
@@
$\qquad=\left[\begin{matrix}0&f_{\rm even}(\frac{g_{1}-g_{2}}{2})&f_{\rm even}(\frac{g_{3}-g_{1}}{2})\\ f_{\rm even}(\frac{g_{1}-g_{2}}{2})&0&f_{\rm even}(\frac{g_{2}-g_{3}}{2})\\ f_{\rm even}(\frac{g_{3}-g_{1}}{2})&f_{\rm even}(\frac{g_{2}-g_{3}}{2})&0\end{matrix}\right]_{ij}\mathbf{X}_{ij}$
for all $i,j\in\{1,2,3\}$. Hence, by summing the last two identities and adding also the part $f_{0}\mathbf{X}$, we arrive at (2.13).
Regarding the eigenvalue multiplicities, the case of a single distinct eigenvalue always trivially leads to (2.11). Next, suppose that $d=3$ and $g_{1}\neq g_{2}=g_{3}$. We proceed just as in (5.4) and (5.5) with the difference that $\mathbf{G}$ and $\mathbf{X}$ are now of the size $3\times 3$. We also note that for the additional elements, we have
@@
for each $i,j\in\{1,2,3\}$, which again leads to the required result. The other two cases $g_{2}\neq g_{3}=g_{1}$ and $g_{3}\neq g_{1}=g_{2}$ are analogous. ∎
@@ 5.2. Symbolic differentiation of matrix functions
$\int_{0}^{1}f^{\prime}\big((1-s)g_{i}+sg_{j}\big)ds=\Big\{\begin{matrix}\frac{f(g_{i})-f(g_{j})}{g_{i}-g_{j}}&g_{i}\neq g_{j}\\ f^{\prime}(g_{i})&\text{else}\end{matrix}\Big\}=\lim_{(x,y)\to(g_{i},g_{j})}\frac{f(x)-f(y)}{x-y}.$
@@
$\frac{df(\mathbf{G})}{d\mathbf{G}}$ $=\int_{0}^{1}f^{\prime}\big((1-s)\operatorname{\mathsf{L}}_{\mathbf{G}}+\,s\operatorname{\mathsf{R}}_{\mathbf{G}}\big)ds\,\mathbf{X}$
$=\mathbf{Q}\Big(\Big[\int_{0}^{1}f^{\prime}\big((1-s)g_{i}+sg_{j}\big)ds\Big]\odot(\mathbf{Q}^{\mathsf{T}}\mathbf{X}\mathbf{Q})\Big)\mathbf{Q}^{\mathsf{T}}$
@@
(B) Using the well-known identity for the beta function ([31, Sect. 6.2])
$\int_{0}^{1}(1-s)^{k}s^{n-1-k}ds=\frac{k!(n-1-k)!}{n!}=\frac{1}{n\binom{n-1}{k}},\quad 0\leq k<n,$
@@
$\frac{df(\mathbf{G})}{d\mathbf{G}}$ $=\int_{0}^{1}n\big((1-s)\operatorname{\mathsf{L}}_{\mathbf{G}}+\,s\operatorname{\mathsf{R}}_{\mathbf{G}}\big)^{n-1}ds$
$=\sum_{k=0}^{n-1}n\binom{n-1}{k}\int_{0}^{1}(1-s)^{k}s^{n-1-k}ds\operatorname{\mathsf{L}}_{\mathbf{G}}^{k}\operatorname{\mathsf{R}}_{\mathbf{G}}^{n-1-k}$
@@
$\frac{de^{\mathbf{G}}}{d\mathbf{G}}=\int_{0}^{1}e^{(1-t)\operatorname{\mathsf{L}}_{\mathbf{G}}+t\operatorname{\mathsf{R}}_{\mathbf{G}}}dt=\int_{0}^{1}e^{(1-t)\operatorname{\mathsf{L}}_{\mathbf{G}}}e^{t\operatorname{\mathsf{R}}_{\mathbf{G}}}dt=\int_{0}^{1}\operatorname{\mathsf{L}}_{e^{(1-t)\mathbf{G}}}\operatorname{\mathsf{R}}_{e^{t\mathbf{G}}}dt.$
@@
$\frac{d\log\mathbf{A}}{d\mathbf{A}}=\frac{\log\operatorname{\mathsf{L}}_{\mathbf{A}}-\log\operatorname{\mathsf{R}}_{\mathbf{A}}}{\operatorname{\mathsf{L}}_{\mathbf{A}}-\operatorname{\mathsf{R}}_{\mathbf{A}}}=\frac{\operatorname{\mathsf{L}}_{\log\mathbf{A}}-\operatorname{\mathsf{R}}_{\log\mathbf{A}}}{e^{\operatorname{\mathsf{L}}_{\log\mathbf{A}}}-e^{\operatorname{\mathsf{R}}_{\log\mathbf{A}}}}$ $=e^{-\operatorname{\mathsf{R}}_{\log\mathbf{A}}}\frac{\operatorname{\mathsf{L}}_{\log\mathbf{A}}-\operatorname{\mathsf{R}}_{\log\mathbf{A}}}{e^{\operatorname{\mathsf{L}}_{\log\mathbf{A}}-\operatorname{\mathsf{R}}_{\log\mathbf{A}}}-1}$
$=\operatorname{\mathsf{R}}_{\mathbf{A}^{-1}}(1/\eta)(2\operatorname{\mathsf{a\!d}}_{\log\mathbf{A}}),$
@@
hence, by integration of this identity over $(0,1)$, we get
@@
$\frac{\sinh(rx)}{\sinh x\cosh((r-1)x)}$ $=\frac{\sinh x\cosh((r-1)x)+\cosh x\sinh((r-1)x)}{\sinh x\cosh((r-1)x)}$
$=1+\frac{\tanh((r-1)x)}{\tanh{x}},$